Database last updated on Wed, 28 Jan 2026 13:40:51 +0000 by localhost
Critical warnings from b-files' comments for days between 2025007 and 2025007 (0 sec, 4 rows)
| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
| rl: | 2025 007 04 27 47 Motor #11: appears stuck with a reference sensor activated. | 4 | 1 | B00725.204 | GOOSE | #204 | 04:21:12 | | 1 | -1 |
| rl: | 2025 007 04 50 27 New value set for watermark LOOP.RUN.TIME: 673.000050 millisec. | 7 | 3 | B00725.204 | GOOSE | #204 | 04:21:20 | | 1 | -1 |
| rl: | 2025 007 04 50 27 New value set for watermark LOOP.RUN.TIME: 673.000050 millisec. | 7 | 3 | B00725.204 | GOOSE | #204 | 04:21:20 | | 1 | -1 |
| rl: | 2025 007 08 34 20 PMT counter failure | 1 | 1 | B00725.204 | GOOSE | #204 | 04:21:40 | | 1 | -1 |
Instrument constant changes for the week before 2025007 (2 sec, 0 rows)
HG burned out? (AP results when HG is ON) (0 sec, 4 rows)
Non-critical warnings from b-files' comments for days between 2025007 and 2025007 (0 sec, 2 rows)
| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
| rl: | 2025 007 04 27 46 Reset Warm -> Tepid from 2025 007 04 27 40 until 2025 007 04 27 41 . Counts: 1,0 | 1 | 3 | B00725.204 | GOOSE | #204 | 04:21:09 | | 1 | -1 |
| rl: | 2025 007 04 27 47 Motor #11: appears stuck with a reference sensor activated. | 1 | 3 | B00725.204 | GOOSE | #204 | 04:21:12 | | 1 | -1 |
Measurements and tests summary.
| _jday | _loc | _serial | _skc | dt | ed | fm | fz | hg | hp | rs | sl | uv | zs |
| 007 | GOOSE BAY | #204 | ecsc,Next schedule will be ecsc (now ecsc ZS= 93 ZS= 93.009) | 2 | 2 | 34 | 20 | 15 | 14 | 2 | 6 | 21 | 38 |
(0 sec)
Relative and absolute humidity for day 2025007 (2 sec, 1 rows)
| Serial | RH, % | AH, g/m3 | 7-day RH history | Des-nt replcd days ago |
| 204 | 2.4 | 0.4 |